# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ 2.2.1. Neighboring Dataset
We first introduce the event-level adjacency ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$. ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ and $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ are called event-level neighboring datasets, denoted as $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$, if $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can be obtained from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by substituting a single record in ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. In this context, the diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$ is defined as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{Event}}}}=\{\mathbf{d}:\|\mathbf{d}\|_{1}\leq 2\}$. This is because when a single record is substituted, $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can deviate from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by 2 entries, with each entry differing by 1, resulting in $\|\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}\|_{1}\leq 2$.
@@
However, it’s important to recognize that users may not be content with protecting only individual actions taken on the internet. Instead, they often seek comprehensive protection for all of their personal information. Consequently, there arises a need for user-level neighboring datasets. To move from event-level neighboring datasets to user-level neighboring datasets, we replace the concept of "the change of one record" with "the change of all records associated with one user." However, it’s important to note that since a user can potentially contribute an infinite number of records, and differential privacy aims to provide worst-case guarantees, introducing excessive noise can significantly impact utility performance. To address this challenge, a common practice is to constrain user contributions, as discussed in prior works such as (dong2023continual, ; googlesql, ; liu2023algorithms, ). This constraint limits each user to having at most $k$ records in the dataset. In cases where a user exceeds this limit and contributes more than $k$ records, only the first $k$ records are retained, while the surplus records are discarded.
@@
2.2.2. Differential Privacy Properties
Theorem 2.6 (Post-processing (zcdp, )).
@@ 2.2.2. Differential Privacy Properties
Theorem 2.7 (Sequential composition (zcdp, )).
@@
Theorem 2.8 (Group privacy for zCDP (zcdp, )).
@@ 3. Related Work
A lack of formal privacy guarantee in industry practice: To address growing concerns over consumer privacy, many advertising networks and platforms have introduced a range of privacy-preserving advertising measurement tools. Notable among these initiatives are Apple’s Private Click Measurement (PCM) framework and Google’s Attribution Reporting API as part of the Privacy Sandbox. PCM mostly relies on reducing data granularity—for example, by allowing only limited user identifiers or implementing delayed reporting—to decrease the amount of information websites can gather about user activities on other sites. Similarly, the Attribution Reporting API delivers event-level reports that share many of PCM’s techniques, such as delayed reporting and coarse data granularity. In addition, the Attribution Reporting API also generates summary reports, or aggregated reports, by utilizing encryption and a trusted execution environment (TEE). Neither PCM nor the Attribution Reporting API can provide any formal privacy guarantee for the advertising measurement output with a mathematically rigorous definition. It is unclear whether such approaches are indeed useful in safeguarding user privacy since any aggregated report could potentially be subject to privacy risks in a way similar to the US census data attack in a series of works by Dwork and others (dwork2017exposed, ; dick2023confidence, ; keller2023database, ). Recently, Google (delaney2024differentially, ) discussed releasing differentially private conversions in a one-shot setting, whereas our focus is on releasing data in an online setting. Alistair (tholoniat2024alistair, ) scheduled the on-device privacy budget efficiently to achieve individual differential privacy (IDP), with a special focus on secure multiparty computation (MPC) and TEE. Additionally, (denison2022private, ) used differentially private stochastic gradient descent to train machine learning models.
@@
Secure computation is not sufficient: Ads industry has recently proposed secure multi-party computation (MPC) based solutions. These solutions ensure the input data encrypted during advertising measurement computations. Examples include PSI-Sum (ion2020deploying, ) by Google and Private Join and Compute (meta, ) by Meta, as well as the Interoperable Private Attribution (IPA) (case2023interoperable, ), a collaborative effort by Mozilla and Meta. Despite the encryption of input data, the reports—output from the MPC computation—are plaintext results without formal privacy guarantees. Recent studies have shown that in the case of advertising measurement, membership inference attacks can exploit this vulnerability, leading to the leakage of user privacy and business confidentiality (guo2022birds, ; Bo, ).
@@
Ensuring privacy guarantees for online reporting poses challenges: An advertising campaign can extend even months. Applying differential privacy to the cumulative ads measurement results ensure the privacy of the final data. However, it is also crucial to report daily measurement results online. Simply applying a differential privacy mechanism by adding independent noise for each day severely degrades measurement accuracy and usability, as demonstrated in the experiments section. One possible research direction to improve the accuracy without sacrificing privacy is to use of tree-based mechanisms to address the online release problem under event-level differential privacy. This research direction was initially introduced by Chan and Dwork (chan2011private, ; dwork2010differential, ). Subsequent studies by Denisov et al. (denisov2022improved, ; henzinger2022constant, ; henzinger2023almost, ) leveraged the matrix mechanism (LMHMR15, ; YZWXYH12, ; li2012adaptive, ; yuan2015optimizing, ; mckenna2018optimizing, ; xiao2020optimizing, ; mckenna2021hdmm, ; xiao2023answering, ) to enhance the accuracy of these tree-based approaches by improving the constant factors in their error rates. (fichtenberger2021differentially, ) studies the continual release problem for graphs.

### v2

@@ 2.2.1 Neighboring Dataset
We first introduce the event-level adjacency ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$. ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ and $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ are called event-level neighboring datasets, denoted as $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}},\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$, if $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can be obtained from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by substituting a single record in ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$. We assume a fixed number of users, and therefore use the replacement notion for defining neighboring datasets. However, with minor adjustments to the sensitivity calculation, our method can also be adapted to the add-or-remove notion. In this context, the diffset of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Event}}}}$ is defined as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{D}^{\text{Event}}}}=\{\mathbf{d}:\|\mathbf{d}\|_{1}\leq 2\}$. This is because when a single record is substituted, $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ can deviate from ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ by 2 entries, with each entry differing by 1, resulting in $\|\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}\|_{1}\leq 2$.
@@
However, it is important to recognize that users may not be content with protecting only individual actions taken on the internet. Instead, they often seek comprehensive protection for all of their personal information. Consequently, there arises a need for user-level neighboring datasets. To move from event-level neighboring datasets to user-level neighboring datasets, we replace the concept of “the change of one record” with “the change of all records associated with one user.” However, it’s important to note that since a user can potentially contribute an infinite number of records, and differential privacy aims to provide worst-case guarantees; introducing excessive noise can significantly impact utility performance. To address this challenge, a common practice is to constrain user contributions, as discussed in prior works such as [19, 49, 33]. This constraint limits each user to having at most $k$ records in the dataset. In cases where a user exceeds this limit and contributes more than $k$ records, only the first $k$ records are retained, while the surplus records are discarded.
@@
2.2.2 Differential Privacy Properties
Theorem 2.7 (Post-processing [8]).
@@ 2.2.2 Differential Privacy Properties
Theorem 2.8 (Sequential composition [8]).
@@
Theorem 2.9 (Group privacy for zCDP [8]).
@@ 3 Related Work
A lack of formal privacy guarantee in industry practices. To address growing consumer concerns over privacy, many advertising networks and platforms have introduced a range of privacy-preserving advertising measurement tools. Notable among these initiatives are Apple’s Private Click Measurement (PCM) framework and Google’s Attribution Reporting API, which is part of the Privacy Sandbox. PCM mostly relies on reducing data granularity—for example, by allowing only limited user identifiers or implementing delayed reporting—to decrease the amount of information websites can gather about user activities on other sites. Similarly, the Attribution Reporting API delivers event-level reports that share many of PCM’s techniques, such as delayed reporting and coarse data granularity. In addition, the Attribution Reporting API also generates summary reports, or aggregated reports, by utilizing encryption and a trusted execution environment (TEE). Neither PCM nor the Attribution Reporting API can provide any formal privacy guarantee for the advertising measurement output with a mathematically rigorous definition. It is unclear whether such approaches are indeed useful in safeguarding user privacy since any aggregated report could potentially be subject to privacy risks in a way similar to the US census data attack in a series of works by Dwork and others [17, 30]. Recently, Google [15] discussed releasing differentially private conversions in a one-shot setting, whereas our focus is on releasing data in an online setting. Alistair [45] scheduled the on-device privacy budget efficiently to achieve individual differential privacy (IDP), with a special focus on secure multiparty computation (MPC) and TEE.
@@
Secure computation is not sufficient. The advertising industry has recently proposed secure multi-party computation (MPC) based solutions. These solutions ensure the input data is encrypted during advertising measurement computations. Examples include PSI-Sum [28] by Google and Private Join and Compute [2] by Meta, as well as the Interoperable Private Attribution (IPA) [10], a collaborative effort by Mozilla and Meta. Despite the encryption of input data, the reports—output from the MPC computation—are plaintext results without formal privacy guarantees. Recent studies have shown that in the case of advertising measurement, membership inference attacks can exploit this vulnerability, potentially risking user privacy and business confidentiality [24, 29].
@@
Ensuring privacy guarantees for online reporting poses challenges. An advertising campaign can extend even months. Applying differential privacy to the cumulative ad measurement results ensure the privacy of the final data. However, it is also crucial to report daily measurement results online. Simply applying a differential privacy mechanism by adding independent noise for each day severely degrades measurement accuracy and usability, as demonstrated by the experiments in Section 6. One possible research direction to improve the accuracy without sacrificing privacy is to use tree-based mechanisms to address the online release problem under event-level differential privacy. This direction was initially introduced by Chan and Dwork [11, 21]. Subsequent studies by Denisov et al. [16] leveraged the matrix mechanism [32, 50, 52] to enhance the accuracy of these tree-based approaches by improving the constant factors in their error rates. However, these methods have limited flexibility, as they primarily focus on minimizing the root mean squared error for prefix-sum queries. In contrast, Ads-BPC can optimize a variety of queries with different objective functions.